# From Complex-Analytic Models to Sparse Domination: A Dyadic Approach of Hypersingular Operators via Bourgain’s Interpolation Method

### v3

@@ Abstract.
In the unit disc setting with $1<t<3/2$, we obtain a full characterization of the $(p,q)$ mapping theory for the dyadic hypersingular maximal operator $\mathcal{M}_{t}^{\mathcal{D}}$, in particular including estimates on the critical line 1/q1/p=2t21/q-1/p=2t-2 and a weighted endpoint criterion in the radial setting. We also prove endpoint estimates for the hypersingular Bergman projection
@@
Our real-variable perspective addresses to an inquiry raised by Cheng–Fang–Wang–Yu on developing effective real-analytic tools in the hypersingular regime for $K_{2t}$, and it also provides a new route toward the critical-line analysis of Forelli–Rudin type operators and related hypersingular operators in both real and complex settings.
@@ 3. $\bullet$
In [12, Section 1], the authors remarked that, in the hypersingular regime $t>1$, developing a workable real-variable approach that yields explicit boundedness results for concrete operators remains an open direction. This is one of the main motivations for the present work.
@@
    Likewise, if one replaces $\mathbb{D}$} by the upper half plane $\mathcal{H}:=\{z\in\mathbb{C}:\textnormal{Im}z>0\}$, then a careful inspection of the argument in (1.2) yields that the operator is again not well-defined on nonzero constant functions whenever $t>1$ (see also [12, Theorem 5] for a related statement concerning the hypersingular Bergman projection $K^{\mathcal{H}}_{2t}$ on the upper half plane).
@@
The results above are summarized in Figure 1.
@@
    (see, [12, Theorem 3] and also Theorem 5.7, (1) for a generalization and strengthening of this result in terms of “hypersingular” sparse operators);
@@
We summarize the above results for $K_{2t}$ in Figure 2.
@@
One major difficulty in analyzing $K_{2t}$ arises at the endpoint $p=\frac{1}{3-2t}$ and $q=1$. Indeed, as observed in [12, Theorem 3], strong-type estimates fail on the critical line 1/q1/p=2t21/q-1/p=2t-2. This is in sharp contrast with the Bergman projection $K_{2}$, whose $L^{2}$ boundedness is simply guaranteed by its definition. As a consequence, both the Calderón–Zygmund approach (for instance, the Calderón–Zygmund decomposition, which uses an $L^{r}$ bound, $1<r<\infty$, to control the bad function) and more recent sparse domination techniques (e.g. the weak-type machinery in [8, Theorem E]) become more delicate to implement directly in this setting.
@@
    The strong-type range above is sharp, as can be seen from the strong-type characterization for $K_{2t}$ (see [12, Theorem 3]). This suggests that the $L^{p}$$L^{q}$ mapping properties of $K_{2t}$ are governed primarily by the underlying geometry of the (dyadic) Carleson boxes, rather than by analyticity.
@@
    From the viewpoint of complex function theory, we advance the study of Forelli–Rudin type operators in the hypersingular regime by establishing new endpoint and critical-line estimates for $K_{2t}$ with $t>1$. To the best of our knowledge, these estimates are new even in the unit disc setting. A key feature of our approach is the introduction of Bourgain’s interpolation method into this setting, which allows us to overcome the lack of strong-type bounds and provides a robust harmonic-analytic framework for hypersingular operators arising in complex variables.
@@
The rest of this paper is organized as follows. In Section 2, we collect basic definitions and notation that will be used throughout the paper. In Section 3, we establish the $L^{p}$ mapping theory for the dyadic hypersingular maximal operator $M_{t}^{\mathcal{D}}$, including both the off-critical-line bounds and the critical-line estimates, together with the endpoint characterizations in the radial weighted setting. In Section 4, we turn to the hypersingular Bergman projection $K_{2t}$ and prove the critical-line estimates, with special emphasis on the endpoint $(p,q)=\left(\frac{1}{3-2t},1\right)$ via a combination of hypersingular sparse domination and Bourgain’s interpolation trick. In Section 5, we introduce hypersingular sparse operators associated with graded sparse families in $\mathbb{R}^{n}$ and prove the corresponding $L^{p}$-$L^{q}$ theory, which in particular extends and strengthens the results obtained earlier for $K_{2t}$. Finally, in Section 6, we propose several open problems and further directions motivated by the methods developed in this paper.
@@ 3. $L^{p}$ theory for the hypersingular maximal operator
3.1. Off-critical line estimate.
@@
Now for any $f\in L^{\frac{1}{3-2t-\varepsilon}}({\mathbb{D}})$, we have to estimate $\left\|{\mathcal{M}}_{t}^{{\mathcal{D}}}f\right\|_{L^{1}({\mathbb{D}})}$. Using (3.1), we have
$\int_{{\mathbb{D}}}{\mathcal{M}}_{t}^{{\mathcal{D}}}f(z)$ $dA(z)=\sum_{k\in\mathbb{Z}}\int_{E_{4^{tk}}\backslash E_{4^{t(k+1)}}}{\mathcal{M}}_{t}^{{\mathcal{D}}}f(z)dA(z)\lesssim\sum_{k\in\mathbb{Z}}4^{tk}\left|E_{4^{tk}}\right|$
$\leq\sum_{k\in\mathbb{Z}}\left(\sum_{i=1}^{\infty}\int_{Q^{\textnormal{up}}_{4^{tk},i}}\left|{\mathcal{M}}^{{\mathcal{D}}}f(z)\right|^{\frac{1}{3-2t-\varepsilon}}dA(z)\right)^{3-2t-\varepsilon}\cdot\left(\sum_{i=1}^{\infty}|Q_{4^{tk},i}|^{\frac{t-1+\varepsilon}{2t+\varepsilon-2}}\right)^{2t+\varepsilon-2}$
(3.3) $\leq\left(\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\int_{Q^{\textnormal{up}}_{4^{tk},i}}\left|{\mathcal{M}}^{{\mathcal{D}}}f(z)\right|^{\frac{1}{3-2t-\varepsilon}}dA(z)\right)^{3-2t-\varepsilon}\cdot\left(\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}|Q_{4^{tk},i}|^{\frac{t-1+\varepsilon}{2t+\varepsilon-2}}\right)^{2t+\varepsilon-2}$
Since for each $k\in\mathbb{Z}$, $\{Q_{4^{tk},i}\}_{i\geq 1}$ is maximal, we have $\left\{Q^{\textnormal{up}}_{4^{tk},i}\right\}_{k\in\mathbb{Z},i\geq 0}$ are mutually disjoint, therefore, the first double sum in (3.1) is bounded above by
@@
$\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}|Q_{4^{tk},i}|^{\frac{t-1+\varepsilon}{2t+\varepsilon-2}}$ $\simeq\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}|Q^{\textnormal{up}}_{4^{tk},i}|^{\frac{t-1+\varepsilon}{2t+\varepsilon-2}}=\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\left|Q^{\textnormal{up}}_{4^{tk},i}\right|\cdot\frac{1}{\left|Q^{\textnormal{up}}_{4^{tk},i}\right|^{\frac{t-1}{2t+\varepsilon-2}}}$
$\simeq\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\int_{Q^{\textnormal{up}}_{4^{tk},i}}\frac{1}{(1-|z|^{2})^{\frac{2(t-1)}{2t+\varepsilon-2}}}dA(z)\lesssim\int_{{\mathbb{D}}}\frac{1}{(1-|z|^{2})^{\frac{2(t-1)}{2t+\varepsilon-2}}}dA(z)$
@@ 3.2. Critical line estimate
The first assertion follows from the off-diagonal Marcinkiewicz interpolation theorem [14, Theorem 1.4.19]. The second follows from by letting $r=\infty$ and the fact that $L^{p}({\mathbb{D}})\subset L^{p,\infty}({\mathbb{D}})$. ∎
@@
An important reason to consider the Békollé–Bonami weights is that they provide a necessary and sufficient condition for $\mathcal{M}$} to extend to a bounded operator on $L^{l}({\mathbb{D}},\omega)$, with operator norm of magnitude $\omega]_{{\bf B}_{l}}^{\frac{1}{l-1}}$ (see, e.g., [1]). Here, $\mathcal{M}$} refers to the standard maximal operator over all Carleson tents.
@@
$\lesssim\left(\sum_{i=1}^{\infty}\int_{Q_{\alpha,i}^{\textnormal{up}}}\left|{\mathcal{M}}^{{\mathcal{D}}}f(z)\right|^{\frac{1}{3-2t}}\omega(z)dA(z)\right)^{3-2t}\left(\sum_{i=1}^{\infty}\int_{Q_{\alpha,i}^{\textnormal{up}}}\frac{dA(z)}{(1-|z|^{2})\omega^{\frac{3-2t}{2t-2}}(z)}\right)^{2t-2}.$
Therefore,
$\int_{{\mathbb{D}}}{\mathcal{M}}_{t}^{{\mathcal{D}}}f(z)$ $dA(z)=\sum_{k\in\mathbb{Z}}\int_{E_{4^{tk}}\backslash E_{4^{t(k+1)}}}{\mathcal{M}}_{t}^{{\mathcal{D}}}f(z)dA(z)\lesssim\sum_{k\in\mathbb{Z}}4^{tk}\left|E_{4^{tk}}\right|$
$\leq\sum_{k\in\mathbb{Z}}\left(\sum_{i=1}^{\infty}\int_{Q_{4^{tk},i}^{\textnormal{up}}}\left|{\mathcal{M}}^{{\mathcal{D}}}f(z)\right|^{\frac{1}{3-2t}}\omega(z)dA(z)\right)^{3-2t}\left(\sum_{i=1}^{\infty}\int_{Q_{4^{tk},i}^{\textnormal{up}}}\frac{dA(z)}{(1-|z|^{2})\omega^{\frac{3-2t}{2t-2}}(z)}\right)^{2t-2}$
$\lesssim\left\|{\mathcal{M}}^{{\mathcal{D}}}f\right\|_{L^{\frac{1}{3-2t}}({\mathbb{D}},\omega)}\cdot\left(\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\int_{Q_{4^{tk},i}^{\textnormal{up}}}\frac{dA(z)}{(1-|z|^{2})\omega^{\frac{3-2t}{2t-2}}(z)}\right)^{2t-2}$
(3.12) $\lesssim[\omega]_{{\bf B}_{\frac{1}{3-2t}}}^{\frac{3-2t}{2t-2}}\cdot\left\|f\right\|_{L^{\frac{1}{3-2t}}({\mathbb{D}},\omega)}\cdot\left(\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\int_{Q_{4^{tk},i}^{\textnormal{up}}}\frac{dA(z)}{(1-|z|^{2})\omega^{\frac{3-2t}{2t-2}}(z)}\right)^{2t-2},$
where in the last estimate, we used the assumption $\omega\in{\bf B}_{\frac{1}{3-2t}}$. Finally for the double summation in (3.2), using the fact that $\{Q_{4^{tk},i}^{\textnormal{up}}\}_{k\in\mathbb{Z},\;i\geq 1}$ are mutually disjoint, we have
$\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\int_{Q_{4^{tk},i}^{\textnormal{up}}}\frac{dA(z)}{(1-|z|^{2})\omega^{\frac{3-2t}{2t-2}}(z)}$ $\lesssim\int_{{\mathbb{D}}}\frac{1}{(1-|z|^{2})\omega^{\frac{3-2t}{2t-2}}(z)}dA(z)$

### v4

@@ Abstract.
In the unit disc setting with $1<t<3/2$, we obtain a full characterization of the $(p,q)$ mapping theory for the dyadic hypersingular maximal operator $\mathcal{M}_{t}^{\mathcal{D}}$, in particular including estimates on the critical line 1/q1/p=2t21/q-1/p=2t-2 and a weighted endpoint criterion in the radial setting. In addition, we establish a novel two-weight estimate for $\mathcal{M}_{t}^{\mathcal{D}}$ in the range $p>q$, valid for any $t>0$. We also prove endpoint estimates for the hypersingular Bergman projection
@@
Our real-variable perspective addresses an inquiry raised by Cheng–Fang–Wang–Yu on developing effective real-analytic tools in the hypersingular regime for $K_{2t}$, and it also provides a new route toward the critical-line analysis of Forelli–Rudin type operators and related hypersingular operators in both real and complex settings.
@@ 3. $\bullet$
In [14, Section 1], the authors remarked that, in the hypersingular regime $t>1$, developing a workable real-variable approach that yields explicit boundedness results for concrete operators remains an open direction. This is one of the main motivations for the present work.
@@
    Likewise, if one replaces $\mathbb{D}$} by the upper half plane $\mathcal{H}:=\{z\in\mathbb{C}:\textnormal{Im}z>0\}$, then a careful inspection of the argument in (1.2) yields that the operator is again not well-defined on nonzero constant functions whenever $t>1$ (see also [14, Theorem 5] for a related statement concerning the hypersingular Bergman projection $K^{\mathcal{H}}_{2t}$ on the upper half plane).
@@
The results above are summarized in Figure 1 below.
@@
To this end, we consider weighted estimates for ${\mathcal{M}}_{t}^{{\mathcal{D}}}$, with the main focus on the regime $\{1\leq p,q\leq\infty:p>q\}$. In this range, we obtain a somewhat more general two-weight estimate for $\mathcal{M}_{t}^{\mathcal{D}}$ for any $t>0$. More precisely, we show that for $\mu,\omega$ being two weights on $\mathbb{D}$} satisfying certain $\bf B_{\infty}$ condition (see, Definition 3.9), ${\mathcal{M}}_{t}^{{\mathcal{D}}}:L^{p}(\omega,{\mathbb{D}})\to L^{q}(\mu,{\mathbb{D}})$ is bounded if and only if $\phi\in L^{\frac{p}{p-q}}({\mathbb{D}})$, where
(1.3) $\phi(z):=\sum_{I\in{\mathcal{D}}}\beta_{I}\mbox{1\hskip-4.25pt{1}}_{Q_{I}^{\textrm{up}}}(z),\quad\textrm{with}\quad\beta_{I}:=\frac{1}{|Q_{I}|^{(t-1)q}}\cdot\frac{\mu(Q_{I})}{|Q_{I}|}\cdot\left(\frac{\sigma(Q_{I})}{|Q_{I}|}\right)^{\frac{q}{p^{\prime}}},$
where $\sigma:=\omega^{-1/(p-1)}$ is the dual weight of $\omega$ (see, Theorem 3.11).
Remark 1.6.
  1. (1)
    The regime $\{1\leq p,q\leq\infty:p\leq q\}$ in the above result can be handled by standard methods in weighted theory (see, e.g., [8] for the case $t=1$).
  2. (2)
    The condition (1.3) may be viewed as a hypersingular counterpart of the Békollé–Bonami condition. Indeed, in the limiting case $p=q$ and $t=1$, (1.3) reduces exactly to the Békollé–Bonami ${\bf B}_{p}$ condition.
  3. (3)
    Condition (1.3) can also be interpreted from the perspective of Bergman–Carleson embeddings in complex function theory, initiated in a series of influential works of Luecking [22, 23, 24]. In that setting, the boundedness of the embedding operator $id:A^{p}({\mathbb{D}})\to L^{q}_{\mu}({\mathbb{D}})$ (here, $A^{p}({\mathbb{D}})$ denotes the standard Bergman space) is characterized by an $L^{\frac{p}{p-q}}({\mathbb{D}})$-integrability condition on an appropriate testing function. From this viewpoint, the above result also suggests a way to understand the Békollé–Bonami condition through Carleson embedding.
@@
    (see, [14, Theorem 3] and also Theorem 5.7, (1) for a generalization and strengthening of this result in terms of “hypersingular” sparse operators);
@@
We summarize the above results for $K_{2t}$ in Figure 2 below.
@@
One major difficulty in analyzing $K_{2t}$ arises at the endpoint $p=\frac{1}{3-2t}$ and $q=1$. Indeed, as observed in [14, Theorem 3], strong-type estimates fail on the critical line 1/q1/p=2t21/q-1/p=2t-2. This is in sharp contrast with the Bergman projection $K_{2}$, whose $L^{2}$ boundedness is simply guaranteed by its definition. As a consequence, both the Calderón–Zygmund approach (for instance, the Calderón–Zygmund decomposition, which uses an $L^{r}$ bound, $1<r<\infty$, to control the bad function) and more recent sparse domination techniques (e.g. the weak-type machinery in [9, Theorem E]) become more delicate to implement directly in this setting.
@@
    The strong-type range above is sharp, as can be seen from the strong-type characterization for $K_{2t}$ (see [14, Theorem 3]). This suggests that the $L^{p}$$L^{q}$ mapping properties of $K_{2t}$ are governed primarily by the underlying geometry of the (dyadic) Carleson boxes, rather than by analyticity.
@@
    From the viewpoint of complex function theory, we develop a systematic theory of the hypersingular maximal operator ${\mathcal{M}}_{t}^{{\mathcal{D}}}$ and also advance the study of Forelli–Rudin type operators in the hypersingular regime by establishing new endpoint and critical-line estimates for $K_{2t}$ with $t>1$. To the best of our knowledge, these estimates are new even in the unit disc setting. A key feature of our approach is the introduction of Bourgain’s interpolation method into this setting, which allows us to overcome the lack of strong-type bounds and provides a robust harmonic-analytic framework for hypersingular operators arising in complex variables.
@@
The rest of this paper is organized as follows. In Section 2, we collect basic definitions and notation that will be used throughout the paper. In Section 3, we establish the $L^{p}$ mapping theory for the dyadic hypersingular maximal operator $M_{t}^{\mathcal{D}}$, including both the off-critical-line bounds and the critical-line estimates, together with the endpoint characterizations in the radial weighted setting. Moreover, we prove a two-weight estimate for ${\mathcal{M}}_{t}^{{\mathcal{D}}},t>0$. In Section 4, we turn to the hypersingular Bergman projection $K_{2t}$ and prove the critical-line estimates, with special emphasis on the endpoint $(p,q)=\left(\frac{1}{3-2t},1\right)$ via a combination of hypersingular sparse domination and Bourgain’s interpolation trick. In Section 5, we introduce hypersingular sparse operators associated with graded sparse families in $\mathbb{R}^{n}$ and prove the corresponding $L^{p}$-$L^{q}$ theory, which in particular extends and strengthens the results obtained earlier for $K_{2t}$. Finally, in Section 6, we propose several open problems and further directions motivated by the methods developed in this paper.
@@ 3. $L^{p}$ theory for the hypersingular maximal operator
3.1. Off-critical line estimate
@@
Now for any $f\in L^{\frac{1}{3-2t-\varepsilon}}({\mathbb{D}})$, we have to estimate $\left\|{\mathcal{M}}_{t}^{{\mathcal{D}}}f\right\|_{L^{1}({\mathbb{D}})}$. Write
[…]
Observe that it can happen that the intersection of the two sets of Carleson boxes $\{Q_{4^{tk},i}\}_{i\geq 1}$ and $\{Q_{4^{t(k+1)},i}\}_{i\geq 1}$ is not empty. Hence, for each $k\in\mathbb{Z}$, define the disjoint union of Carleson boxes
[…]
Note that
  1. $\bullet$
    $E_{4^{tk}}\backslash E_{4^{t(k+1)}}\subseteq\bigcup_{i\geq 1}\widetilde{Q}_{4^{tk},i}$.
  2. $\bullet$
    $\{\widetilde{Q}^{\textrm{up}}_{4^{tk},i}\}_{k\geq\mathbb{Z},\;i\geq 1}$ are mutually disjoint.
Using now (3.1) with $E_{4^{tk}}$ replaced by $E_{4^{tk}}\backslash E_{4^{t(k+1)}}$, and $\{Q_{4^{tk},i}\}_{i\geq 1}$ replaced by $\{\widetilde{Q}_{4^{tk},i}\}_{i\geq 1}$, respectively, we have
RHS of (3.3) $\leq\sum_{k\in\mathbb{Z}}\left(\sum_{i=1}^{\infty}\int_{\widetilde{Q}^{\textnormal{up}}_{4^{tk},i}}\left|{\mathcal{M}}^{{\mathcal{D}}}f(z)\right|^{\frac{1}{3-2t-\varepsilon}}dA(z)\right)^{3-2t-\varepsilon}\cdot\left(\sum_{i=1}^{\infty}|\widetilde{Q}_{4^{tk},i}|^{\frac{t-1+\varepsilon}{2t+\varepsilon-2}}\right)^{2t+\varepsilon-2}$
(3.5) $\leq\left(\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\int_{\widetilde{Q}^{\textnormal{up}}_{4^{tk},i}}\left|{\mathcal{M}}^{{\mathcal{D}}}f(z)\right|^{\frac{1}{3-2t-\varepsilon}}dA(z)\right)^{3-2t-\varepsilon}\cdot\left(\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}|\widetilde{Q}_{4^{tk},i}|^{\frac{t-1+\varepsilon}{2t+\varepsilon-2}}\right)^{2t+\varepsilon-2}$
Since $\{\widetilde{Q}^{\textnormal{up}}_{4^{tk},i}\}_{k\in\mathbb{Z},\;i\geq 1}$ are mutually disjoint, therefore, the first double sum in (3.1) is bounded above by
@@ 2. $\bullet$
$\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}|\widetilde{Q}_{4^{tk},i}|^{\frac{t-1+\varepsilon}{2t+\varepsilon-2}}$ $\simeq\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}|\widetilde{Q}^{\textnormal{up}}_{4^{tk},i}|^{\frac{t-1+\varepsilon}{2t+\varepsilon-2}}=\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\left|\widetilde{Q}^{\textnormal{up}}_{4^{tk},i}\right|\cdot\frac{1}{\left|\widetilde{Q}^{\textnormal{up}}_{4^{tk},i}\right|^{\frac{t-1}{2t+\varepsilon-2}}}$
$\simeq\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\int_{\widetilde{Q}^{\textnormal{up}}_{4^{tk},i}}\frac{1}{(1-|z|^{2})^{\frac{2(t-1)}{2t+\varepsilon-2}}}dA(z)\lesssim\int_{{\mathbb{D}}}\frac{1}{(1-|z|^{2})^{\frac{2(t-1)}{2t+\varepsilon-2}}}dA(z)$
@@ 3.2. Critical line estimate
The first assertion follows from the off-diagonal Marcinkiewicz interpolation theorem [16, Theorem 1.4.19]. The second follows from by letting $r=\infty$ and the fact that $L^{p}({\mathbb{D}})\subset L^{p,\infty}({\mathbb{D}})$. ∎
@@
An important reason to consider the Békollé–Bonami weights is that they provide a necessary and sufficient condition for
  1. $\bullet$
    the standard Hardy–Littlewood maximal operator over all Carleson tents $\mathcal{M}$} to extend to a bounded operator on $L^{l}({\mathbb{D}},\omega)$, with operator norm of magnitude $\omega]_{{\bf B}_{l}}^{\frac{1}{l-1}}$ (see, e.g., [1, 30]).
  2. $\bullet$
    the Bergman projection $\mathcal{P}$ acting as a bounded operator on $L^{l}({\mathbb{D}},\omega)$ (see, e.g., [5, 30, 31]).
@@ 2. $\bullet$
(3.14) $\lesssim\left(\sum_{i=1}^{\infty}\int_{Q_{\alpha,i}^{\textnormal{up}}}\left|{\mathcal{M}}^{{\mathcal{D}}}f(z)\right|^{\frac{1}{3-2t}}\omega(z)dA(z)\right)^{3-2t}\left(\sum_{i=1}^{\infty}\int_{Q_{\alpha,i}^{\textnormal{up}}}\frac{dA(z)}{(1-|z|^{2})\omega^{\frac{3-2t}{2t-2}}(z)}\right)^{2t-2}.$
To estimate $\left\|{\mathcal{M}}_{t}^{{\mathcal{D}}}f\right\|_{L^{1}({\mathbb{D}})}$, write
[…]
Similarly as in (3.4), let
[…]
Then using (3.2) with again $E_{4^{tk}}$ replaced by $E_{4^{tk}}\backslash E_{4^{t(k+1)}}$, and $\{Q_{4^{tk},i}\}_{i\geq 1}$ replaced by $Q_{4^{tk},i}\}_{i\geq 1}\backslash\{Q_{4^{t(k+1)},i}\}_{i\geq 1}$, respectively, we have
RHS of (3.15)
$\leq\sum_{k\in\mathbb{Z}}\left(\sum_{i=1}^{\infty}\int_{\widetilde{Q}_{4^{tk},i}^{\textnormal{up}}}\left|{\mathcal{M}}^{{\mathcal{D}}}f(z)\right|^{\frac{1}{3-2t}}\omega(z)dA(z)\right)^{3-2t}\left(\sum_{i=1}^{\infty}\int_{\widetilde{Q}_{4^{tk},i}^{\textnormal{up}}}\frac{dA(z)}{(1-|z|^{2})\omega^{\frac{3-2t}{2t-2}}(z)}\right)^{2t-2}$
$\lesssim\left\|{\mathcal{M}}^{{\mathcal{D}}}f\right\|_{L^{\frac{1}{3-2t}}({\mathbb{D}},\;\omega)}\cdot\left(\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\int_{\widetilde{Q}_{4^{tk},i}^{\textnormal{up}}}\frac{dA(z)}{(1-|z|^{2})\omega^{\frac{3-2t}{2t-2}}(z)}\right)^{2t-2}$
(3.16) $\lesssim[\omega]_{{\bf B}_{\frac{1}{3-2t}}}^{\frac{3-2t}{2t-2}}\cdot\left\|f\right\|_{L^{\frac{1}{3-2t}}({\mathbb{D}},\;\omega)}\cdot\left(\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\int_{\widetilde{Q}_{4^{tk},i}^{\textnormal{up}}}\frac{dA(z)}{(1-|z|^{2})\omega^{\frac{3-2t}{2t-2}}(z)}\right)^{2t-2},$
where in the last estimate, we used the assumption $\omega\in{\bf B}_{\frac{1}{3-2t}}$. Finally for the double summation in (3.2), using the fact that $\{\widetilde{Q}_{4^{tk},i}^{\textnormal{up}}\}_{k\in\mathbb{Z},\;i\geq 1}$ are mutually disjoint, we have
$\sum_{k\in\mathbb{Z}}\sum_{i=1}^{\infty}\int_{\widetilde{Q}_{4^{tk},i}^{\textnormal{up}}}\frac{dA(z)}{(1-|z|^{2})\omega^{\frac{3-2t}{2t-2}}(z)}$ $\lesssim\int_{{\mathbb{D}}}\frac{1}{(1-|z|^{2})\omega^{\frac{3-2t}{2t-2}}(z)}dA(z)$